### v1

# FourierKAN-GCF: Fourier Kolmogorov-Arnold Network - An Effective and Efficient Feature Transformation for Graph Collaborative Filtering

Abstract.
Graph Collaborative Filtering (GCF) has achieved state-of-the-art performance for recommendation tasks. However, most GCF structures simplify the feature transformation and nonlinear operation during message passing in the graph convolution network (GCN). We revisit these two components and discover that a part of feature transformation and nonlinear operation during message passing in GCN can improve the representation of GCF, but increase the difficulty of training.
In this work, we propose a simple and effective graph-based recommendation model called FourierKAN-GCF. Specifically, it utilizes a novel Fourier Kolmogorov-Arnold Network (KAN) to replace the multilayer perceptron (MLP) as a part of the feature transformation during message passing in GCN, which improves the representation power of GCF and is easy to train. We further employ message dropout and node dropout strategies to improve the representation power and robustness of the model. Extensive experiments on two public datasets demonstrate the superiority of FourierKAN-GCF over most state-of-the-art methods. The implementation code is available at https://github.com/Jinfeng-Xu/FKAN-GCF.
Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Transform
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems;
1. Introduction
Recommender system has been widely used to alleviate information overload on the web (Xu et al., 2024). The aim of recommender system is to predict whether a user will interact with an unobserved item. To this end, collaborative filtering (CF) addresses it by learning user preferences from similar users. Many works (He et al., 2020; Wang et al., 2019; Jiang et al., 2023) harnessing graph-based recommendation to capture latent user preferences from user-item interactions have gained notable successes. The main reason is that the interactive data between users and items are naturally graph-structured. For example, NGCF (Wang et al., 2019) utilizes whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation. However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary for feature extraction in recommendation field and further proposes a light GCN. Most of the subsequent graph-based models (Zhou et al., 2023a; Wei et al., 2019; Yu et al., 2023) are based on LightGCN for further exploration.
However, there are many unfair analyses of NGCF in LightGCN: a) There are two different feature transformations in NGCF, but LightGCN removed them without performing separate fine-grained ablation experiments. b) LightGCN also removed the interaction information representation part of NGCF without giving reasons. Therefore, we question: ‘Is feature transformation during message passing in GCN really unnecessary in recommendation?’
We provide an empirical analysis for NGCF and LightGCN in Section 2. Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but the training process becomes more complex. In this work, we introduce a simple yet powerful graph-based recommendation model called FourierKAN-GCF. Particularly, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs. This substitution enhances the representational capabilities and training ease of the GCF. Additionally, we implement various dropout techniques to further boost the model’s representational strength and robustness. Extensive experiments on public datasets demonstrate the superiority of our FourierKAN-GCF over state-of-the-art methods.
2. Preliminary
2.1. NGCF Brief
NGCF retains feature transformation and nonlinear operation during the message passing in GCN, formally:
(1) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
$\mathbf{e}_{i}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are two trainable weight matrices to perform feature transformation in each layer. $\sigma$ is the nonlinear activation function. It is worth noting that $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ can be regarded as the aggregated representation from neighbors, while $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ and $\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})$ can be regarded as the aggregated representation from interaction information. The final embeddings are calculated as:
(2) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)},$
where ||||| | is the concatenation operation. NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-00, since ego layer-00 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
2.2. LightGCN Brief
LightGCN provides an analysis of the feature transformation and nonlinear operation in NGCF. It mainly provides four observations:
1. Removing the entire feature transformation including both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ leads to consistent improvements over NGCF.
2. Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance.
3. Removing both entire feature transformation and nonlinear operation can improve performance significantly.
4. The deterioration of NGCF stems from the training difficulty rather than over-fitting.
To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, LightGCN leverages the user-item graph to propagate embeddings as:
(3) $\mathbf{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},\quad\mathbf{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},$
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. LightGCN only retains the aggregation of neighbors’ representation but removes the aggregation of interaction information representation. The final embeddings are calculated as:
(4) $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{\mathbf{e}_{u}^{(l)}}{L+1},\quad\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{\mathbf{e}_{i}^{(l)}}{L+1},$
where $L$ is the total layer number. LightGCN considers the ego layer-00 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ in NGCF.
We argue that observation 4 is the main reason why LightGCN performs better than NGCF in most cases. Observations 1-3 were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction. Besides, we argue that LightGCN verified the feature transformation not contributing to learning better features by comparing NGCF with removed both $\mathbf{W}_{1}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$, which is not a fair comparison. We point out that $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ naturally contain information about user preferences and item properties that can be adequately described through multiple feature dimensions. Therefore, $\mathbf{W}_{1}$ is an unnecessary feature transformation part. However, we point out that the interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ contains valuable interaction information, which can not be easily extracted by the heuristic rule. In this case, $\mathbf{W}_{2}$ is a necessary feature transformation part that contributes to feature extraction. In the next subsection, we provide empirical evidence for these arguments.
Table 1. Performance comparison of LightGCN, NGCF, and four variants of NGCF in terms of Recall@20 and NDCG@20.
Dataset Metrics LightGCN NGCF NGCF-f1 NGCF-f2 NGCF-i NGCF-n
MOOC Recall@20 0.3307 0.3361 0.3377$\uparrow$ 0.3357$\downarrow$ 0.3301$\downarrow$ 0.3343$\downarrow$
NDCG@20 0.1811 0.1894 0.1926$\uparrow$ 0.1897$\uparrow$ 0.1824$\downarrow$ 0.1878$\downarrow$
Games Recall@20 0.0447 0.0379 0.0414$\uparrow$ 0.0388$\uparrow$ 0.0362$\downarrow$ 0.0373$\downarrow$
NDCG@20 0.0227 0.0196 0.0209$\uparrow$ 0.0192$\downarrow$ 0.0181$\downarrow$ 0.0196--
2.3. Re-Analysis for NGCF
We implement four simplified variants of NGCF:
  • NGCF-f1, which removes feature transformation matrix $\mathbf{W}_{1}$.
  • NGCF-f2, which removes feature transformation matrix $\mathbf{W}_{2}$.
  • NGCF-i, which removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$.
  • NGCF-n, which removes the nonlinear operation $\sigma$.
We keep all hyper-parameter settings the same as the optimal settings of NGCF. As Table 1 shows, we conclude the findings that:
1. For both MOOC and Games datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements. Therefore, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary.
2. For MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration. However, for Games dataset, it will lead to a small improvement. We own this to that user-item interaction in MOOC dataset has higher density than in Games dataset.
3. Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets. It shows that this part can bring valuable interaction information.
4. Nonlinear operation only brings a small positive effect.
5. NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult.
Then, we can draw some conclusions. First, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary for NGCF. Besides, the feature transformation matrix $\mathbf{W}_{2}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ is beneficial for feature extraction. Last, nonlinear operation occupies only a minor influence.
3. Methodology
In this section, we detail the overall architecture of our FourierKAN-GCF, NGCF, and LightGCN in Figure 1. Our Fourier Kolmogorov-Arnold Network provides a powerful interaction representation aggregation part in NGCF to achieve a significant performance improvement and make it easy to train.
Figure 1. Overview of FourierKAN-GCF, NGCF, and LightGCN
3.1. Kolmogorov-Arnold Network (KAN)
Kolmogorov-Arnold Network (Liu et al., 2024) is a promising alternative to Multi-Layer Perceptron (MLP). MLP is inspired by the universal approximation theorem (Cybenko, 1989). KAN focuses on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1961). KAN can be formed as:
(5) $\operatorname{KAN}=f(\mathbf{x})=\sum_{q=1}^{2n+1}\Phi_{q}(\sum_{p=1}^{n}\phi_{q,p}(\mathbf{x}_{p})),$
where $\phi_{q,p}$ are univariate functions that map each input variable $\mathbf{x}_{p}$ such $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. $\phi_{q,p}(\mathbf{x}_{p})$ is trainable activation function. In the KolmogovArnold theorem, the inner functions form a KAN layer with $n_{in}$ = $n$ and $n_{out}$ = $12n+$ 1, and the outer functions form a KAN layer with $n_{in}$ = $12n+$ 1 and $n_{out}$ = $n$. So the Kolmogorov-Arnold representations in this formula are simply compositions of two KAN layers. A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\phi(\mathbf{x})$ is the sum of the basis function $b(\mathbf{x})$ and the spline function:
(6) $\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x})).$
$\operatorname{spline}(\mathbf{x})$ is parameterized as a linear combination of B-splines:
(7) $b(x)=\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}},\quad\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x}),$
where $c_{i}$ is trainable.
3.2. Fourier KAN
However, KAN is more difficult to train than MLP due to the spline function, which does not satisfy our motivation to provide a more effective and efficient method to replace $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ part. The essence of KAN is to form an arbitrary function by superposition of multiple nonlinear functions. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier transform (Nussbaumer and Nussbaumer, 1982) is an interesting choice. Therefore, we propose the following equation:
(8) $\phi_{F}(\mathbf{x})=\sum_{i=1}^{d}\sum_{k=1}^{g}\left(\cos\left(k\mathbf{x}_{i}\right)\cdot a_{ik}+\sin\left(k\mathbf{x}_{i}\right)\cdot b_{ik}\right),$
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier transform corresponding to each input dimension. The Fourier transform has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
3.3. FourierKAN-GCF
The message passing in our FourierKAN-GCF is defined as:
(9) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}+\phi_{F}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
$\mathbf{e}_{i}^{(l+1)}$ $=\sigma(\mathbf{e}_{i}^{(l)}+\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}+\phi_{F}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
where $\phi_{F}(\cdot)$ is simplified single layer Fourier KAN function. We remove the unnecessary transform matrix $\mathbf{W}_{1}$ in NGCF and utilize our Fourier KAN function to replace the transform matrix $\mathbf{W}_{2}$. The final user embedding and item embedding are calculated as:
(10) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{i}^{(L)},$
where ||||| | is the concatenation operation.
3.4. Dropout Strategies
To further prevent the over-fitting problem for FourierKAN-GCF, we adopt message dropout and node dropout strategies. For the message dropout, we randomly change some message passing being propagated in Eq 9 by zero, with a probability $1-p_{m}$, where $p_{m}$ is the message dropout ratio. For the node dropout, we randomly drop $1-p_{n}$ nodes of the matrix, where $p_{n}$ is the node dropout ratio.
Table 2. Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K. The superscript indicates the improvement is statistically significant where the p-value is less than 0.05.
Datasets 1MOOC 2Games
Metrics R@10 R@20 R@50 N@10 N@20 N@50 R@10 R@20 R@50 N@10 N@20 N@50
BPR-MF 0.2401 0.3353 0.4813 0.1571 0.1898 0.2261 0.0210 0.0369 0.0699 0.0135 0.0183 0.0265
BUIR 0.2354 0.3196 0.4801 0.1589 0.1835 0.2224 0.0227 0.0384 0.0749 0.0143 0.0192 0.0282
NGCF 0.2453 0.3361 0.4799 0.1722 0.1894 0.2349 0.0231 0.0379 0.0782 0.0142 0.0196 0.0274
LR-GCCF 0.2466 0.3336 0.4809 0.1678 0.1938 0.2294 0.0253 0.0440 0.0815 0.0171 0.0224 0.0317
LightGCN 0.2402 0.3307 0.4773 0.1581 0.1811 0.2217 0.0271 0.0447 0.0844 0.0174 0.0227 0.0326
UltraGCN 0.2372 0.3194 0.4701 0.1737 0.1962 0.2307 0.0277 0.0459 0.0844 0.0173 0.0230 0.0331
IMP-GCN 0.2170 0.2788 0.4183 0.1531 0.1717 0.2057 0.0273 0.0461 0.0839 0.0171 0.0232 0.0323
w/o MD 0.2518 0.3523 0.4912 0.1814 0.2116 0.2449 0.0267 0.0452 0.0825 0.0165 0.0221 0.0314
w/o ND 0.2603 0.3527 0.4839 0.1800 0.2071 0.2439 0.0257 0.0452 0.0809 0.0161 0.0219 0.0309
KAN-GCF 0.2536 0.3417 0.4984 0.1758 0.2024 0.2396 0.0272 0.0451 0.0837 0.0172 0.0229 0.0325
FourierKAN-GCF 0.2595 0.3564 0.5065 0.1866 0.2147 0.2462 0.0287 0.0473 0.0856 0.0176 0.0252 0.0342
1Dataset can be accessed at http://moocdata.cn. 2Dataset can be accessed at http://jmcauley.ucsd.edu/data/amazon/links.html.
3.5. Model Training
To learn model parameters, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss function as our optimization criterion. The core objective of BPR is to enhance the divergence in the predictive preference between positive and negative items within each user-item triplet $(u,i_{p},i_{n})\in\mathcal{D}$, where $\mathcal{D}$ signifies the collection of training data, the term positive item $p$ pertains to an item with which the user $u$ has interacted, whereas the negative item $n$ is selected randomly from the pool of items without interaction with user $u$.
(11) $\mathcal{L}=\sum_{(u,i_{p},i_{n})\in\mathcal{D}}-\ln\sigma(\mathbf{\hat{e}}_{u}^{T}\mathbf{\hat{e}}_{i_{p}}-\mathbf{\hat{e}}_{u}^{T}\mathbf{\hat{e}}_{i_{n}})+\lambda\|\mathbf{\Theta}\|^{2},$
where $\lambda$ controls the $L_{2}$ regularization strength, $\sigma$ is the Sigmoid function, and $\mathbf{\Theta}$ denotes model parameters.
Table 3. Statistics of experimental datasets.
Dataset # Users # Items # Interaction Sparsity
MOOC 82535 1302 458453 99.57%
Games 50677 16897 454529 99.95%
4. Experiment
4.1. Datasets and Evaluation Metrics
We conduct our experiments on two real-world datasets: MOOC and Anazon Video Games (Games for short). Details can be found in Table 3. For a fair comparison, we follow previous work (Zhou et al., 2023b) to sort all the observed user-item interactions of each dataset in chronological order based on the interaction timestamps. Then, we split each dataset with a ratio of 7:1:2 for training, validation, and testing. Regarding evaluation metrics, we adopt two well-established metrics: Recall@K (R@K) and NDCG@K (N@K). We report the average metrics of all users in the test dataset among $K$ = 10, $K$ = 20, and $K$ = 50.
4.2. Baselines and Experimental Settings
To verify the effectiveness of FourierKAN-GCF, we select two traditional matrix factorization models (BPR-MF (Rendle et al., 2012) and BUIR (Lee et al., 2021)) and five state-of-the-art GCN-based models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) for comparison. For a fair comparison, we implement all baselines under a unified framework (Zhou et al., 2023b), fix the embedding size of both users and items to 64 for all models, initialize the embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2014) as the optimizer. Besides, we carefully tune the hyper-parameters of each baseline following their published papers to find optimal settings. For our FourierKAN-GCF, we tune the $\lambda$ of $L_{2}$ regularization term in {0, 1e-2, 1e-1, 1, 10}, and tune the layer number $L$ from 1 to 4. The gridsize $g$ of Fourier KAN is searched from {1, 2, 4, 8}. We further grid search the message dropout ratio $p_{m}$ and node dropout ratio $p_{n}$ in {0.0, 0.1, 0.2, 0.3}.
4.3. Performance Comparison
The results of our experiments are listed in Table 2. We can identify that the majority of GCN-based methods outperform traditional matrix factorization methods, which is a common trend. Our FourierKAN-GCF outperforms all baselines in two datasets across various metrics. We owe our superiority to FourierKAN, which is easier to train and has greater representation power than MLP. It is worth noting that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
4.4. Ablation Study
Table 2 also demonstrates the significance of dropout strategies. We use w/o MD and w/o ND to denote without message dropout and without node dropout, respectively. This ablation study shows that both message dropout and node dropout play a distinct role in improving model representation power and model robustness. We use KAN-GCF to denote a variant that utilizes our standard KAN function to replace transform matrix $\mathbf{W}_{2}$ in NGCF. It shows that standard KAN is better than MLP but worse than our Fourier KAN.
4.5. Sensitivity Analysis
To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on two datasets with different hyper-parameters. The optimal layer number $L$ is 3 on all datasets, shown in Figure LABEL:fig:l. Besides, Figure LABEL:fig:g demonstrates that for MOOC and Games datasets, optimal ($p_{m}$, $p_{n}$) pairs are (0.2, 0.2) and (0.1, 0.1), respectively. As Figure LABEL:fig:dropout illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size. For the relatively sparse dataset Games, $g$ = 2 is the best grid size. This further demonstrates that FourierKAN can be easily adapted to various datasets.
5. Conclusion
In this paper, we revisit the processes of feature transformation and nonlinear operations during message passing in the NGCF. We assert that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN but add complexity in training. To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN. Inspired by KAN, FourierKAN employs a Fourier transform instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power. Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies. Extensive experiments on public datasets demonstrate the superiority of our model over the state-of-the-art methods.