### v3

# Enhancing Graph Collaborative Filtering with FourierKAN Feature Transformation

Abstract.
Graph Collaborative Filtering (GCF) has emerged as a dominant paradigm in modern recommendation systems, excelling at modeling complex user-item interactions and capturing high-order collaborative signals through graph-structured learning. Most existing GCF models predominantly rely on simplified graph architectures like LightGCN, which strategically remove feature transformation and activation functions from vanilla graph convolution networks. Through systematic analysis, we reveal that feature transformation in message propagation can enhance model representation, though at the cost of increased training difficulty. To this end, we propose FourierKAN-GCF, a novel GCN framework that adopts Fourier Kolmogorov-Arnold Networks as efficient transformation modules within graph propagation layers. This design enhances model representation while decreasing training difficulty. Our FourierKAN-GCF can achieve higher recommendation performance than most widely used GCF backbone models. In addition, it can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance. Extensive experiments on three public datasets demonstrate the superiority of FourierKAN-GCF.
Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Coefficients
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3760909isbn: 979-8-4007-2040-6/2025/11
1. Introduction
Recommender systems are widely used to alleviate information overload on the Web (Chen et al., 2020; He et al., 2017; Chen et al., 2025a, b; Xu et al., 2025e, a), aiming to recommend suitable items for users based on their historical behavior. Collaborative filtering (CF) addresses this by learning user preferences from similar users. Recently, graph-based CF (GCF) models (He et al., 2020; Wang et al., 2019; Jiang et al., 2023; Xu et al., 2025d, c, b, 2024b) have achieved notable success, as user-item interactions are naturally graph-structured. For example, NGCF (Wang et al., 2019) adopts whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation. However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary in the recommendation field and further proposes a lightweight GCN. Consequently, most subsequent GCF models (Chen et al., 2020; Cai et al., 2023; Xu et al., 2024a) adopt LightGCN as the backbone for further exploration.
However, there are many unfair analyses of NGCF in LightGCN: a) There are two different feature transformations in NGCF, but LightGCN removed them without performing separate fine-grained ablation experiments. b) LightGCN also removed the interaction information representation part of NGCF without giving reasons. Therefore, we question: ‘Is feature transformation during message passing in GCN really unnecessary in recommendation?’
We provide an empirical analysis for NGCF and LightGCN in Section 2. Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but increases the training difficulty. In this work, we introduce a simple yet powerful graph-based recommendation model called FourierKAN-GCF. Specifically, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs. This substitution enhances the representational capabilities and reduces the difficulty of training for GCFs. FourierKAN-GCF can achieve higher recommendation performance than most widely-used GCF backbone models. In addition, FourierKAN-GCF can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance. Extensive experiments on public datasets demonstrate the superiority of FourierKAN-GCF over state-of-the-art methods. Our work is intended to reawaken researchers’ thinking about feature transformation in GCF, rather than arbitrarily removing it.
2. Preliminary
2.1. NGCF Brief
NGCF (Wang et al., 2019) retains feature transformation and nonlinear operation during the message passing in GCN, formally:
(1) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
$\mathbf{e}_{i}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are two trainable weight matrices to perform feature transformation in each layer. $\sigma$ is the nonlinear activation function. It is worth noting that $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ can be regarded as the aggregated representation from neighbors, while $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ and $\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})$ can be regarded as the aggregated representation from interaction information. The final embeddings are calculated by $\mathbf{\hat{e}}_{u}=\{\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)}\}$ and $\mathbf{\hat{e}}_{i}=\{\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)}\}$, where $||$ is the concatenation operation. NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-0, since ego layer-0 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
2.2. LightGCN Brief
LightGCN (He et al., 2020) analyzes the feature transformation and nonlinear operation in NGCF. It offers four observations: a) Removing the entire feature transformation, including $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, leads to consistent improvements over NGCF. b) Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance. c) Removing both entire feature transformation and nonlinear operation can improve performance significantly. d) The deterioration of NGCF stems from the training difficulty rather than over-fitting.
To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, the user-item graph to propagate embeddings as:
(2) $\mathbf{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},\quad\mathbf{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},$
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. LightGCN only retains the aggregation of neighbors’ representation but removes the aggregation of interaction information representation. The final embeddings are calculated as $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{1}{L+1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{1}{L+1}\mathbf{e}_{i}^{(l)}$, where $L$ is the total layer number. LightGCN considers the ego layer-0 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$.
We argue that observation (d) is the main reason why LightGCN performs better than NGCF in most cases. Observations (a-c) were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction. Besides, we argue that LightGCN verified the feature transformation not contributing to learning better features by comparing NGCF with removed both $\mathbf{W}_{1}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$, which is not a fair comparison. We point out that $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ naturally contain information about user preferences and item properties that can be adequately described through multiple feature dimensions. Therefore, $\mathbf{W}_{1}$ is an unnecessary feature transformation part. However, we point out that the interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ contains valuable interaction information, which can not be easily extracted by the heuristic rule. In this case, $\mathbf{W}_{2}$ is a necessary feature transformation part that contributes to feature extraction. In the next subsection, we provide an empirical analysis.
Table 1. Performance comparison of LightGCN, NGCF, and six variants of NGCF in terms of Recall@20 (R@20) and NDCG@20 (N@20).
Datasets MOOC Amazon Gowalla
Metrics R@20 N@20 R@20 N@20 R@20 N@20
LightGCN 0.3307 0.1811 0.0447 0.0227 0.1830 0.1152
NGCF 0.3361 0.1894 0.0379 0.0196 0.1755 0.1013
[…]
2.3. Re-Analysis for NGCF
We implement six variants of NGCF: 1) NGCF-f1 removes feature transformation matrix $\mathbf{W}_{1}$. 2) NGCF-f2 removes feature transformation matrix $\mathbf{W}_{2}$. 3) NGCF-i removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$. 4) NGCF-f1-f2 removes both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. 5) NGCF-f1-i removes both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$. 6) NGCF-n removes the nonlinear operation $\sigma$.
We keep all optimal hyper-parameter settings as the NGCF reported. As Table 1 shows, we conclude the findings: a) For all three datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements. Therefore, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary. b) For the MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration. However, for the other two datasets, it will lead to a small improvement. Moreover, removing both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ will lead to a small deterioration than only removing $\mathbf{W}_{1}$ for all datasets. We own this phenomenon to that removing only $\mathbf{W}_{1}$ can make the model optimization focus on training $\mathbf{W}_{2}$, and remove both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ will lose the representation power from feature transformation. c) Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets. Additionally, removing both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will make a obviously degration than only remove $\mathbf{W}_{1}$ and even NGCF. This shows that this part can bring valuable information about interaction. d) Nonlinear operation only brings a small positive effect. e) NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult (He et al., 2020). Then, we can draw some conclusions. First, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary for NGCF. Besides, the feature transformation matrix $\mathbf{W}_{2}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ is beneficial for feature extraction. Last, nonlinear operation only has a minor influence.
To this end, a feature transformation approach, simpler to train than MLPs but with strong representational capabilities, could be effective in graph learning for recommendations. KAN (Liu et al., 2024) is regarded as a promising alternative to MLP. While KAN shares the same theoretically unlimited representational capacity as MLP (Wang et al., 2024), the practical representational capacity of MLP is constrained by the hidden dimensionality. In contrast, the practical representational capacity of KAN depends on its ability to fit trainable activation functions. Our work can be seen as exploring adopting KAN as a feature transformation approach within GCF.
3. Methodology
We detail the overview of FourierKAN-GCF [^1] Code is available at: https://anonymous.4open.science/r/FourierKAN-GCF-r. in Figure 1. FourierKAN-GCF introduces a promising feature transformation for GCF, which significantly boosts performance and simplifies training. Note that FourierKAN-GCF can be adopted as the backbone in existing advanced self-supervised models.
Figure 1. Overview of FourierKAN-GCF.
3.1. Kolmogorov-Arnold Network (KAN)
Kolmogorov-Arnold Network (Liu et al., 2024) is a promising alternative to Multi-Layer Perceptron (MLP). MLP is inspired by the universal approximation theorem (Cybenko, 1989). KAN focuses on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1961). Specifically, unlike MLPs, which have fixed activation functions on nodes, KANs contain learnable activation functions on edges (weights). This unique architecture enables KANs to learn nonlinear functions more effectively. Formally:
(3) $\operatorname{KAN}=f(\mathbf{x})=\sum_{q=1}^{2n+1}\Phi_{q}(\sum_{p=1}^{n}\phi_{q,p}(\mathbf{x}_{p})),$
where $\phi_{q,p}$ are univariate functions that map each input variable $\mathbf{x}_{p}$ such $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. $\phi_{q,p}(\mathbf{x}_{p})$ is trainable activation function. In the KolmogovArnold theorem, the inner functions form a KAN layer with $_{in} = nn$ and $n_{out}$ = $2n+1$, and the outer functions form a KAN layer with $n_{in}$ = $2n+1$ and $_{out} = nn$. So the Kolmogorov-Arnold representations in this formula are simply compositions of two KAN layers. A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}}$. $\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x}))$ is the sum of the basis function $b(\mathbf{x})$ and function $\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x})$ is a linear combination of B-splines, where $c_{i}$ is a trainable parameter.
3.2. Fourier KAN
To further reduce the training difficulty and be able to adopt to different scenarios. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is a potential choice. Therefore, we propose the following equation:
(4) $\phi_{F}(\mathbf{x})=\sum_{i=1}^{d}\sum_{k=1}^{g}\left(\cos\left(k\mathbf{x}_{i}\right)\cdot a_{ik}+\sin\left(k\mathbf{x}_{i}\right)\cdot b_{ik}\right),$
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier Coefficients corresponding to each input dimension. The Fourier Coefficients has a significant advantage in computational efficiency and reduces the training difficulty caused by the spline function.
3.3. FourierKAN-GCF
The message passing in FourierKAN-GCF is defined as:
(5) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}+\phi_{F}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
$\mathbf{e}_{i}^{(l+1)}$ $=\sigma(\mathbf{e}_{i}^{(l)}+\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}+\phi_{F}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
where $\phi_{F}(\cdot)$ is simplified single layer Fourier KAN function. We remove the unnecessary transform matrix $\mathbf{W}_{1}$ in NGCF and utilize our Fourier KAN function to replace the transform matrix $\mathbf{W}_{2}$. The final user embedding and item embedding are calculated by $\mathbf{\hat{e}}_{u}=\{\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)}\}$ and $\mathbf{\hat{e}}_{i}=\{\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)}\}$, where $||$ is the concatenation operation.
3.4. Dropout Strategies
To mitigate overfitting in FourierKAN-GCF, we employ message dropout and node dropout strategies, similar to NGCF (Wang et al., 2019). In message dropout, a fraction $1-p_{m}$ of message passing in Eq. 5 is randomly set to zero, where $p_{m}$ is the dropout ratio. In node dropout, $1-p_{n}$ of the nodes in the matrix are randomly dropped, with $p_{n}$ as the dropout ratio.
3.5. Model Training
For model optimization, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2009) loss function as our optimization criterion. The core objective of BPR is to enhance the divergence in the predictive preference between positive and negative items within each user-item triplet $(u,i_{p},i_{n})\in\mathcal{D}$, where $\mathcal{D}$ signifies the collection of training data, the term positive item $p$ pertains to an item with which the user $u$ has interacted, whereas the negative item $n$ is selected randomly from the pool of items without interaction with user $u$.
(6) $\mathcal{L}=\sum_{(u,i_{p},i_{n})\in\mathcal{D}}-\ln\sigma(\mathbf{\hat{e}}_{u}^{T}\mathbf{\hat{e}}_{i_{p}}-\mathbf{\hat{e}}_{u}^{T}\mathbf{\hat{e}}_{i_{n}})+\lambda\|\mathbf{\Theta}\|^{2},$
where $\lambda$ controls the $L_{2}$ regularization strength, $\sigma$ is the Sigmoid function, and $\mathbf{\Theta}$ denotes model parameters.
4. Experiment
In this section, we compare the performance of FourierKAN-GCF with popular graph-based backbone models and demonstrate its compatibility with advanced self-supervised graph-based models.
Table 2. Statistics of experimental datasets.
Dataset # Users # Items # Interaction Sparsity
MOOC 82,535 1,302 458,453 99.57%
Amazon 50,677 16,897 454,529 99.95%
Gowalla 29,859 40,989 1,027,464 99.92%
4.1. Datasets and Evaluation Metrics
Table 3. Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K. The superscript indicates the improvement is statistically significant where the p-value is less than 0.05. #T denotes seconds per epoch.
Datasets MOOC Amazon Gowalla
Metrics R@20 R@50 N@20 N@50 #T R@20 R@50 N@20 N@50 #T R@20 R@50 N@20 N@50 #T
NGCF 0.3361 0.4799 0.1894 0.2349 5.2s 0.0379 0.0782 0.0196 0.0274 4.4s 0.1755 0.2811 0.1013 0.1270 37.5s
LR-GCCF 0.3336 0.4809 0.1938 0.2294 5.1s 0.0440 0.0815 0.0224 0.0317 4.7s 0.1803 0.2971 0.1101 0.1369 40.1s
LightGCN 0.3307 0.4773 0.1811 0.2217 4.3s 0.0447 0.0844 0.0227 0.0326 3.5s 0.1830 0.3044 0.1152 0.1414 26.4s
UltraGCN 0.3194 0.4701 0.1962 0.2307 4.9s 0.0459 0.0844 0.0230 0.0331 4.0s 0.1798 0.2909 0.1059 0.1328 33.8s
IMP-GCN 0.2788 0.4183 0.1717 0.2057 54.2s 0.0461 0.0839 0.0232 0.0323 44.9s 0.1808 0.2932 0.1060 0.1345 143.6s
KAN-GCF 0.3417 0.4984 0.2024 0.2396 4.9s 0.0451 0.0837 0.0229 0.0325 4.0s 0.1922 0.3023 0.1142 0.1403 32.9s
FourierKAN-GCF 0.3564 0.5065 0.2147 0.2462 4.6s 0.0473 0.0856 0.0252 0.0342 3.7s 0.1962 0.3077 0.1179 0.1436 29.8s
w/o MD 0.3523 0.4912 0.2116 0.2449 4.5s 0.0452 0.0825 0.0221 0.0314 3.7s 0.1920 0.3031 0.1150 0.1408 29.5s
w/o ND 0.3527 0.4839 0.2071 0.2439 4.5s 0.0452 0.0809 0.0219 0.0309 3.6s 0.1908 0.3017 0.1133 0.1389 29.6s
Table 4. Performance comparison of different backbones on advanced recommendation models in terms of R@K and N@K.
Datasets MOOC Amazon Gowalla
Metrics Backbones R@20 R@50 N@20 N@50 R@20 R@50 N@20 N@50 R@20 R@50 N@20 N@50
SimGCL LightGCN 0.3503 0.5032 0.2109 0.2428 0.0462 0.0837 0.0230 0.0327 0.2028 0.3126 0.1184 0.1459
FourierKAN-GCF 0.3639 0.5188 0.2216 0.2543 0.0482 0.0868 0.0258 0.0352 0.2084 0.3163 0.1234 0.1521
LightGCL LightGCN 0.2742 0.4139 0.1697 0.2018 0.0453 0.0818 0.0225 0.0318 0.1888 0.2976 0.1081 0.1348
FourierKAN-GCF 0.2823 0.4291 0.1733 0.2072 0.0479 0.0864 0.0250 0.0347 0.2007 0.3100 0.1198 0.1452
RecDCL LightGCN 0.3531 0.5009 0.2113 0.2423 0.0469 0.0848 0.0235 0.0334 0.1993 0.3052 0.1160 0.1433
FourierKAN-GCF 0.3608 0.5120 0.2175 0.2514 0.0484 0.0863 0.0253 0.0349 0.2051 0.3113 0.1203 0.1472
We conduct experiments on three real-world datasets: MOOC, Amazon Video Games (Amazon), and Gowalla. Details can be found in Table 2. For a fair comparison, we sort all observed user-item interactions chronologically based on the interaction timestamps. Then, we split each dataset with a ratio of 7:1:2 for training, validation, and testing. Regarding evaluation metrics, we adopt two well-established metrics (Järvelin and Kekäläinen, 2002): Recall@K (R@K) and NDCG@K (N@K).
4.2. Baselines and Experimental Settings
To verify the effectiveness of FourierKAN-GCF, we select five GCN-based backbone models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) as baselines. Moreover, FourierKAN-GCF can also be a more powerful backbone to enhance existing advanced self-supervised enhanced graph-based models (SimGCL (Yu et al., 2022), LightGCL (Cai et al., 2023), and RecDCL (Zhang et al., 2024)). For a fair comparison, we fix the embedding size of both users and items to 64 for all models, initialize embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2015) as optimizer. Besides, we tune the hyper-parameters of each baseline following their published papers. For FourierKAN-GCF, we fix the $\lambda$ = $1$ for $L_{2}$ regularization, and tune layer number $L$ from 1 to 4. The gridsize $g$ is searched from {1, 2, 4, 8}. Message and node dropout ratios $p_{m}$ and $p_{n}$ is searched from {0.0, 0.1, 0.2, 0.3}.
4.3. Performance Comparison
The results of our experiments are listed in Table 3. Our FourierKAN-GCF outperforms all baselines in three datasets across various metrics. Moreover, FourierKAN-GCF is second only to LightGCN in efficiency, which demonstrates it effectively reduces the training difficulty associated with feature transformation. We owe our superiority to FourierKAN, which is easier to train and has greater representation power than MLP. Note that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
4.4. Compatibility Analysis
We further evaluate the compatibility of FourierKAN-GCF with advanced self-supervised graph-based models by replacing their backbones. As shown in Table 4, adopting FourierKAN-GCF significantly improves performance, highlighting the importance of feature transformation in GCF and establishing FourierKAN-GCF as an effective solution.
4.5. Ablation Study
Table 3 also demonstrates the significance of dropout strategies. We use w/o MD and w/o ND to denote without message dropout and without node dropout, respectively. This ablation study shows that both message dropout and node dropout play a distinct role in improving model representation power and model robustness. We use KAN-GCF to denote a variant that utilizes our standard KAN function to replace transform matrix $\mathbf{W}_{2}$ in NGCF. It shows that standard KAN is better than MLP but worse than our Fourier KAN.
(a) Layer number $L$ (Recall@20)
(b) Grid size $g$ (Recall@20)
Figure 2. Effective of layer number $L$ and grid size $g$.
(a) MOOC
(b) Amazon
(c) Gowalla
Figure 3. Study of dropout ratio pair $p_{m}$ and $p_{n}$ (Recall@20).
4.6. Sensitivity Analysis
To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on three datasets with different hyper-parameters. The optimal layer number $L$ is 3 on all datasets, shown in Figure 2(a). Besides, Figure 2(b) demonstrates that for MOOC dataset, the optimal ($p_{m}$, $p_{n}$) pair is (0.2, 0.2), and for Amazon and Gowalla datasets, the optimal pairs are all (0.1, 0.1). As Figure 3 illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size. For the relatively sparse datasets Amazon and Gowalla, $g$ = 2 is the best grid size. This further demonstrates that FourierKAN can be easily adapted to various datasets.
5. Conclusion
In this paper, we revisit feature transformation and nonlinear operations in the message-passing mechanism of GCNs. While feature transformation enhances interaction representation and boosts performance, it increases training complexity. To this end, we introduce a new feature transformation for GCF called FourierKAN. Inspired by KAN, FourierKAN employs the Fourier Coefficients instead of the Spline function in standard KAN. We further propose a simple yet effective GCF model (FourierKAN-GCF), which reduces the difficulty of training. In addition, FourierKAN-GCF can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance. Extensive experiments on public datasets verify the superiority of our model over the advanced methods.
GenAI Usage Disclosure
No GenAI tools were used in any stage of the research, nor in the writing.
Acknowledgements.
This work was supported by the Hong Kong UGC General Research Fund no. 17203320 and 17209822, and the project grants from the HKU-SCF FinTech Academy.